OpenStudy (anonymous):

What are the units of k in the following expression: rate = k [A][B]? A) s-1 B) M-1s-1 C) M-2s-1 D) M

OpenStudy (anonymous):

Will give medal!!

OpenStudy (anonymous):

@dan815 @iGreen

OpenStudy (jfraser):

\(rate\) always has to have units of mol/(L*s), so the units of k must be changed to match based on the rate order.

OpenStudy (anonymous):

How do you do that?

OpenStudy (anonymous):

@JFraser

OpenStudy (jfraser):

the equation (with units) must look like this

OpenStudy (jfraser):

\[\frac{mol}{L*s} = k*\frac{mol}{L}*\frac{mol}{L}\]

OpenStudy (jfraser):

what do the units of \(k\) have to be in order to make that a true statement?